# SelectTTS: Synthesizing Anyone’s Voice via Discrete Unit-Based Frame Selection

### v1

@@ Abstract
Synthesizing the voices of unseen speakers is a persisting challenge in multi-speaker text-to-speech (TTS). Most multi-speaker TTS models rely on modeling speaker characteristics through speaker conditioning during training. Modeling unseen speaker attributes through this approach has necessitated an increase in model complexity, which makes it challenging to reproduce results and improve upon them. We design a simple alternative to this. We propose SelectTTS, a novel method to select the appropriate frames from the target speaker and decode using frame-level self-supervised learning (SSL) features. We show that this approach can effectively capture speaker characteristics for unseen speakers, and achieves comparable results to other multi-speaker TTS frameworks in both objective and subjective metrics. With SelectTTS, we show that frame selection from the target speaker’s speech is a direct way to achieve generalization in unseen speakers with low model complexity. We achieve better speaker similarity performance than SOTA baselines XTTS-v2 and VALL-E with over an 8x reduction in model parameters and a 270x reduction in training data.
@@ I Introduction
Recent text-to-speech approaches [1, 2, 3] have demonstrated that given sufficient data and large enough model capacity, TTS models are capable of producing speech of remarkably high quality and naturalness. While the development of large-scale TTS models [4] offers its benefits, it also introduces challenges associated with model reproducibility. This issue becomes especially relevant in multi-speaker TTS frameworks.
Multi-speaker TTS for unseen speakers is a challenging problem as the objective is multifold - learning text-to-speech semantic prediction while simultaneously capturing the speaker timbre and acoustics[5]. Previously, multi-speaker TTS predominantly relied on speaker labels and embeddings [6, 7] to condition models to learn speaker traits [8, 9]. Speaker embeddings were initially created for tasks such as speaker recognition [10], but applying these embeddings to model unseen speaker characteristics in synthesis has proven to be an uphill task [11]. More recently, zero-shot TTS frameworks have started to exploit neural codec language modeling [2, 12]. They replace traditional mel-spectrograms with audio codec codes and use in-context learning capability to enable prompt-based zero-shot TTS.
Both the embedding and language modeling approach rely on speaker conditioning, either in the form of speaker embeddings or acoustic prompts, requiring the network to learn the modeling of speaker timbre given the condition. While this has led to state-of-the-art (SOTA) performance [13, 14, 15], with the cost of increased model capacity and complexity, as well as data requirements [13, 16], hindering reproducibility and further development of those methods. With SelectTTS, we propose a straightforward and effective alternative to conditional learning of speaker traits where speaker modeling is performed by non-parametric frame selection of SSL features. This approach allows training a more parameter efficient and less complex model without huge data requirements, while still achieving performance comparable to larger models in reproducing speaker timbre for multi-speaker TTS.
This is enabled by the advancements in self-supervised learning speech models [17, 18]. These SSL features show tremendous potential in capturing the linguistic, speaker, and prosody information [19]. Their masked prediction objective [20] ensures that they possess strong semantic information while also being able to capture speaker acoustic information, which is necessary for multi-speaker TTS. A recent work, kNN-VC [21], has directly leveraged this capability of SSL features in unit selection-based voice conversion [22] by replacing each frame-level feature in the source utterance with the closest neighbours in the reference speech to construct the target feature sequence. It demonstrates that selecting SSL features on a frame-by-frame basis from the reference speech can synthesize voices with a high degree of similarity to the target speaker. Persuaded by the idea of selecting frames in the SSL feature space, we develop frame selection algorithms to leverage the generalization ability of SSL for synthesizing voices of unseen speakers through multi-speaker TTS.
@@
With SelectTTS, we introduce a new paradigm of frame selection-based multi-speaker TTS that directly utilizes frames from unseen target speakers to clone their voice. Rather than conditioning input text with speaker representations or using speech tokens as prompts to model speaker timbre, our method splits the task into two stages. In the first stage, we predict frame-level semantic units from text to model speech semantic content. In the second stage, we obtain intermediate representations with the necessary speaker and acoustic information by selecting frame-level features from the reference speech according to the predicted semantic units. Combining frame selection with rich SSL features achieves SOTA speaker similarity, as it uses the target speaker’s frames directly. This framework, with its multi-stage structure and non-parametric approach, offers reduced model complexity and greater customization compared to larger baselines. The main contributions of the paper are summarized as: 1) We propose a multi-speaker text-to-speech strategy that completely separates and simplifies the tasks of semantic prediction and speaker modeling in TTS, making the overall framework easily reproducible and open to further development; 2) We introduce novel frame selection algorithms, sub-sequence matching and inverse k-means sampling, that directly select frames from the target speaker to accurately reproduce speaker timbre; 3) We utilize the discrete SSL features for text-to-semantic modeling and frame selection; the continuous SSL features for vocoding, showing the benefits of leveraging both feature spaces.
@@ II Related work
Learning speaker characteristics has been the primary goal in multi-speaker TTS research. In the deep learning era, traditional methods are mostly based on conditioning the neural networks with speaker representations[23, 24, 3] or pre-trained speaker embeddings[7, 25]. On the other hand, inspired by the success of LLMs, in-context learning with large codec-based models, such as VALL-E[2], VoiceCraft[13] and Voicebox[1], leverages bi-directional context for speech-infilling tasks using speaker information in the form of acoustic prompts, achieving SOTA speaker similarity with the cost of training in very large-scale data. Methods such as SPEAR-TTS[26], UniCATS [27] reduce the data requirement compared to those large models while using acoustic prompt conditioning, however, they remain to be complex models. Additionally, methods like UnitSpeech[28] and HierSpeech[29] perform speaker adaptation through fine-tuning. In this work, we explore a new alternative approach and show that a simple non-parametric frame selection method can achieve SOTA speaker similarity results.
@@ III SelectTTS
We propose SelectTTS, a multi-speaker TTS framework that directly utilizes frames from the unseen speaker for decoding speech. The proposed method includes two training stages and an offline intermediate stage. In the first training stage, we train a text-to-semantic-unit model. We perform frame selection using the predicted semantic units to choose the frames from the target speaker. Frame selection is performed in an offline manner and helps us recover the continuous SSL feature from its discrete counterpart, as illustrated in Fig. 1(b). Finally, we train the vocoder to map these frame-level continuous SSL features to speech waveform. In this section, we describe in detail about each one of these stages.
@@ III-A Semantic unit tokenizers
SelectTTS relies on its frame selection pipeline to choose the correct frames from the reference speech. We perform frame selection in the discrete semantic-unit space. We use the following tokenizers to bridge the modality gap between text and speech.
@@ III-A1 Speech-unit-tokenizer
The speech-unit tokenizer converts speech into frame-level sequences of semantic units. Continuous SSL features are extracted from the speech at the frame level using a pre-trained SSL model. These features are then converted into discrete semantic units through k-means clustering, as illustrated in Fig. 1(a). We denote this transformation as $z$ = SpeechUnitTokenizer($Z$), where $z$ represents the discrete unit sequence and $Z$ is the continuous SSL feature sequence.
@@ III-A2 Text-unit-tokenizer
We train a non-auto-regressive model that learns to predict frame-level discrete semantic units from text. For this task, we use an architecture that is based on FastSpeech2 [30]. This model consists of a text encoder, a duration prediction model, and a unit decoder. This tokenizer is trained on a parallel dataset containing text and speech discrete units. The speech discrete units are obtained by using the SpeechUnitTokenizer transform. We use an external alignment tool [31] to obtain phoneme durations which is used as the target to train the duration predictor. At inference, this tokenizer consumes non-aligned phoneme sequences and predicts frame-level speech discrete units. The duration model predicts the phoneme durations and accordingly upsamples the phonemes. This tokenizer is denoted by TextUnitTokenizer. We emphasize that predicting discrete units instead of continuous features simplifies the task and reduces overall complexity, allowing for more effective frame selection algorithms. With discrete units, we can easily select sequences of frames together, rather than individually as with continuous features which would be limited for capturing intra-frame dependencies.
@@ III-B1 Sub-sequence matching
We formulate the sub-sequence matching problem as follows: Let $\hat{z}$ = TextUnitTokenizer(text) represent a sequence of predicted semantic units obtained from the TextUnitTokenizer. Similarly, let $z_{\text{ref}}$ = SpeechUnitTokenizer(speech) denote the semantic unit sequence derived from the reference speech of the target speaker using the speech-unit tokenizer. In our sub-sequence matching algorithm, we find sub-sequences of the predicted sequence $\hat{z}$ that is present in the reference unit sequence $z_{\text{ref}}$. Once we find these sub-sequences, we replace the sub-sequence-matched discrete unit sequence with the corresponding continuous SSL features from the reference speech as shown in Fig. 1(b). We start from the longest sub-sequence length and iteratively replace the discrete units till there are no more sub-sequences. In our sub-sequence matching implementation, we begin by searching for matches with a maximum length of 10 to maintain computational efficiency, with a minimum length set at 2. The intuition behind sub-sequence matching is that - by choosing chunks of speech segments at a time, we hope to get the most accurate match in the form of real speech segments instead of relying on only frame-level selection to reduce artifacts and improve segment-level prosody.
@@ III-B2 Inverse k-means sampling
For frames that do not have sub-sequence matches, we replace the discrete unit with a continuous SSL feature belonging to the same discrete unit cluster. This method is referred to as inverse k-means sampling because it performs the reverse operation of k-means clustering, recovering the continuous SSL features from the discretized SSL feature space. We also notice that the discrete unit cluster may be an empty set in the case when the reference speech is short. In this case, we replace the discrete unit with an SSL feature from the nearest k-means cluster that is not empty. For SSL feature sampling, we experimented with two different approaches: randomly sampling one of the frames from the target unit cluster or using the average of all SSL features from the target unit cluster in the reference speech.
@@ III-C Vocoder
The vocoder is used to convert the continuous WavLM features into an audio waveform. The vocoder can be trained without any dependence on frame selection algorithms as we can compute the frame-level features of the ground truth audio and feed that as the input to vocoding. However, this causes a mismatch between the inputs during training and inference. To overcome this mismatch, we perform frame selection with the ground truth audio before training the vocoder. Given an input utterance in a set of utterances from a speaker, we choose the reference speech to be every other utterance from the same speaker. We extract ground truth WavLM units using SpeechUnitTokenizer from the input utterance and perform frame selection described in section III-B. The selected frame-level SSL features are then used as input to the vocoder. The target for the vocoder is the ground-truth audio. In our preliminary listening experiments, we noticed that fine-tuning the vocoder with frames selected with sub-sequence matching and inverse k-means sampling significantly reduces artifacts.
@@ IV-A SelectTTS implementation
The choice of SSL feature is an important decision in our work. Since we are training a text-to-semantic-unit model, initially we experimented with layers 22 and 23 of WavLM-Large [18] since these layers achieve the best performance on the phone recognition task and have the most content information. However, our preliminary synthesis experiments suggested that this causes poor reproduction of speaker and prosody [32], which is essential in multi-speaker TTS. Finally, we use layer 6 on WavLM-large, like kNN-VC, since it has a high degree of speaker information as indicated by its speaker identification performance.
@@ IV-A1 Speech-unit-tokenizer
The WavLM-Large encoder produces continuous feature vectors for every 20ms of 16kHz of audio. We discretize this using k-means as described in Section III-A1. We use 2000 cluster centers for our proposed method after initially experimenting with 100, 500 clusters as more fine-grained units resulted in better intelligibility.
@@ IV-A2 Text-unit-tokenizer
We modify the original FastSpeech2 [30] architecture to design a text-to-semantic unit predictor. We replace the prediction of spectrum with the prediction of semantic units. The ground truth units are derived by applying the SpeechTokenizer on the ground truth audio. The network is optimized using the Adam optimizer. We use a learning rate of 5e-4 and a batch size of 10k phonemes. On an Nvidia RTX 3090 GPU, the network reaches convergence within 20 minutes or after approximately 10,000 steps. We build on top of SpeechLM’s [33] implementation for this tokenizer.
IV-B Vocoder
We use a HiFi-GAN V1 architecture from [34]. We start with the pre-trained vocoder model from the kNN-VC implementation [^1] https://github.com/bshall/knn-vc. We fine-tune two vocoders with frames selected by proposed algorithms: one where frame selection is based solely on inverse k-means sampling, and another where frame selection incorporates both inverse k-means sampling and sub-sequence matching.

### v2

@@ Abstract
Synthesizing the voices of unseen speakers remains a persisting challenge in multi-speaker text-to-speech (TTS). Existing methods model speaker characteristics through speaker conditioning during training, leading to increased model complexity and limiting reproducibility and accessibility. A lower-complexity method would enable speech synthesis research with limited computational and data resources to reach to a wider use. To this end, we propose SelectTTS, a simple and effective alternative. SelectTTS selects appropriate frames from the target speaker and decodes them using frame-level self-supervised learning (SSL) features. We demonstrate that this approach can effectively capture speaker characteristics for unseen speakers and achieves performance comparable to state-of-the-art multi-speaker TTS frameworks on both objective and subjective metrics. By directly selecting frames from the target speaker’s speech, SelectTTS enables generalization to unseen speakers with significantly lower model complexity. Compared to baselines such as XTTS-v2 and VALL-E, SelectTTS achieves better speaker similarity while reducing model parameters by over 8× and training data requirements by 270×.
@@ I Introduction
Recent text-to-speech approaches [1, 2, 3] have demonstrated that given sufficient data and large enough model capacity, TTS models are capable of producing speech of remarkably high quality and naturalness. While the development of large-scale TTS models [4] brings significant advantages, it also introduces challenges associated with model reproducibility. This issue becomes especially relevant in multi-speaker TTS frameworks. Multi-speaker TTS for unseen speakers is a challenging problem as the objective is multifold - learning text-to-speech semantic prediction while simultaneously capturing the speaker timbre and acoustics[5]. Previously, multi-speaker TTS predominantly relied on speaker labels and embeddings [6, 7] to condition models to learn speaker traits [8, 9]. More recently, zero-shot TTS frameworks have advanced this area by leveraging neural codec language modeling [2, 10] combined with acoustic prompting for speaker timbre control.
Both the embedding and language modeling approach rely on speaker conditioning, either in the form of speaker embeddings or acoustic prompts, requiring the network to learn the conditional modeling of speaker timbre. While leading to state-of-the-art (SOTA) performance [11, 12, 13], it comes at the cost of increased model complexity, capacity, and data requirements [11, 14], limiting reproducibility and further development. In particular, these models are computationally expensive to train and deploy, and their reliance on large-scale computational and data resources makes them impractical for many research and application settings. To address these limitations, we propose SelectTTS, a simple and effective alternative to conditional learning of speaker traits where speaker modeling is performed by non-parametric frame selection of SSL features. This approach enables training more parameter-efficient and less complex models without extensive data requirements, while still achieving speaker timbre reproduction comparable to larger models.
This is made possible by SSL speech models [15, 16], which have demonstrated strong capabilities in capturing linguistic, speaker, and prosodic information [17]. The masked prediction objectives used in these models [18] encourage the learning of rich semantic representations while preserving speaker identity. A recent work, kNN-VC [19], has leveraged this property by performing unit selection-based voice conversion, replacing each frame-level feature in the source utterance with its nearest neighbors in the reference speech. This frame-by-frame selection of SSL features enables the synthesis of voices with high similarity to the target speaker.
@@
With SelectTTS, we introduce a new paradigm of frame selection-based multi-speaker TTS that directly utilizes frames from unseen target speakers to clone their voice. Instead of conditioning the input text with speaker embeddings or using speech tokens as prompts to model speaker timbre, SelectTTS splits the task into two stages. In the first stage, frame-level semantic units are predicted from text to model the speech content. In the second stage, intermediate representations enriched with speaker and acoustic information are obtained by selecting frame-level features from the reference speech based on the predicted semantic units. Combining frame selection algorithms with rich SSL features achieves SOTA speaker similarity, as it directly leverages frames from the target speaker. This multi-stage, non-parametric framework reduces model complexity and offers greater customization compared to larger baseline systems.
The main contributions of this work are as follows: 1) We propose a multi-speaker TTS strategy that completely separates and simplifies the tasks of semantic prediction and speaker modeling in TTS, making the overall framework easily reproducible and open to further development; 2) We introduce novel frame selection algorithms, sub-sequence matching and inverse k-means sampling, that directly select frames from the target speaker to accurately reproduce speaker timbre; 3) We utilize the discrete SSL features for text-to-semantic prediction and frame selection; the continuous SSL features for vocoding, showing the benefits of leveraging both feature spaces.
@@ II Related work
Learning speaker characteristics has been a primary focus of multi-speaker TTS research. In the deep learning era, traditional approaches have primarily relied on conditioning neural networks with speaker representations [20, 21, 3, 7, 22]. More recently, inspired by large language models (LLMs), in-context learning with large codec-based models, such as VALL-E[2], VoiceCraft[11], Voicebox[1] have been explored. These models leverage bi-directional context for speech-infilling tasks using speaker information in the form of acoustic prompts, achieving SOTA speaker similarity with the cost of training in very large-scale data. Methods such as SPEAR-TTS[23], UniCATS [24] reduce the data requirement compared to those models while using acoustic prompting, yet, they remain to be complex. Additionally, methods such as UnitSpeech[25], HierSpeech[26] perform speaker adaptation by fine-tuning. In this work, we explore an alternative direction and show that a simple, non-parametric frame selection method can achieve SOTA speaker similarity without the complexity of large-scale models.
@@ III SelectTTS
SelectTTS consists of two training stages and one offline intermediate stage. In the first training stage, we train a text-to-semantic-unit prediction model. Using the predicted semantic units, we perform frame selection by identifying corresponding frames from the target speaker. This offline frame selection step enables the recovery of continuous SSL features from their discrete counterparts, as illustrated in Fig. 1(b). Finally, in the second training stage, we train a vocoder to map the selected frame-level continuous SSL features to the speech waveform. In the following subsections, we describe each stage of the framework in detail.
@@ III-A Semantic unit tokenizers
SelectTTS relies on its frame selection pipeline to choose the correct frames from the reference speech. To enable effective matching, frame selection is performed in a discrete semantic-unit space.
@@ III-A1 Speech-unit-tokenizer
The speech-unit tokenizer converts speech into frame-level sequences of semantic units. Continuous SSL features are extracted from the speech at the frame level using a pre-trained SSL model. These features are then converted into discrete semantic units through k-means clustering, as in Fig. 1(a). We denote this transformation as $z$ = SpeechUnitTokenizer($Z$), where $z$ represents the discrete unit sequence and $Z$ is the continuous SSL feature sequence.
@@ III-A2 Text-unit-tokenizer
We train a non-autoregressive model to predict frame-level discrete semantic units directly from text. For this task, we adopt an architecture based on FastSpeech2 [27], consisting of a text encoder, a duration prediction module, and a unit decoder. The model is trained on a parallel dataset containing text and corresponding discrete speech units, where the speech units are obtained using the SpeechUnitTokenizer transform. An external alignment tool [28] is used to obtain phoneme durations, which serve as supervision for training the duration predictor.
During inference, the TextUnitTokenizer consumes non-aligned phoneme sequences and predicts frame-level sequences of discrete semantic units. The duration model predicts phoneme durations and upsamples the phoneme sequence accordingly to match the frame-level granularity. We emphasize that predicting discrete units instead of continuous features simplifies the modeling task and reduces overall system complexity. Moreover, operating in the discrete space enables more effective frame selection, allowing sequences of frames to be selected together, thereby capturing intra-frame dependencies that would be difficult to handle with continuous features.
@@ III-B1 Sub-sequence matching
We formulate the sub-sequence matching problem as follows: Let $\hat{z}$ = TextUnitTokenizer(text) represent a sequence of predicted semantic units obtained from the TextUnitTokenizer. Similarly, let $z_{\text{ref}}$ = SpeechUnitTokenizer(speech) denote the semantic unit sequence derived from the reference speech of the target speaker using the speech-unit tokenizer. In our sub-sequence matching algorithm, we find sub-sequences of the predicted sequence $\hat{z}$ that is present in the reference unit sequence $z_{\text{ref}}$. Once we find these sub-sequences, we replace the sub-sequence-matched discrete unit sequence with the corresponding continuous SSL features from the reference speech as in Fig. 1(b). We start from the longest sub-sequence length and iteratively replace the discrete units till there are no more matches. In the implementation, we begin by searching for matches with a maximum length of 10 to maintain computational efficiency, with a minimum length of 2. The intuition of sub-sequence matching is that by choosing chunks of speech segments at a time, we hope to get the most accurate match in the form of real speech segments instead of relying on only frame-level selection, to reduce artifacts and improve segment-level prosody.
@@ III-B2 Inverse k-means sampling
For frames without direct sub-sequence matches, we apply a strategy called inverse $k$-means sampling. In this approach, we identify frames from the reference speech that belong to the same discrete unit cluster as the predicted semantic unit. This effectively performs the reverse of the $k$-means clustering process by retrieving continuous SSL features from the discretized semantic space. If the predicted unit cluster is not represented in the reference speech - typically due to the limited reference duration - we instead select frames from the nearest nonempty cluster. Once the appropriate cluster is determined, the corresponding continuous SSL feature is obtained either by randomly sampling a frame from the target cluster or by averaging all SSL features associated with that cluster in the reference speech.
@@ III-C Vocoder
The vocoder converts continuous WavLM features into audio waveforms. It can be trained with ground-truth frame-level feature and speech waveform pairs, independently from frame selection algorithms. However, this causes a mismatch between the inputs during training and inference. To overcome this mismatch, we perform frame selection with the ground truth audio before training the vocoder. Given an input utterance in a set of utterances from a speaker, we choose the reference speech to be every other utterance from the same speaker. We extract ground truth WavLM units using SpeechUnitTokenizer from the input utterance and perform frame selection described in section III-B. The selected frame-level SSL features are then used as input to the vocoder. The target for the vocoder is the ground-truth audio. In preliminary experiments, we observe that fine-tuning the vocoder using subsequence-matched features significantly reduces synthesis artifacts and improves audio quality.
III-D SelectTTS vs kNN-VC
SelectTTS is inspired by the retrieval-based VC method kNN-VC [19], but it introduces a fundamentally different task and framework. Unlike SSL feature-based VC, where SSL features can be directly extracted from input speech, SelectTTS predicts semantic representations from text and performs frame selection from reference speech to synthesize new utterances. This design bridges text and speaker-specific SSL features, enabling efficient, low-complexity TTS for unseen speakers.
@@ IV-A SelectTTS implementation
The choice of SSL feature layer is critical in our framework. We initially experimented with layers 22 and 23 of WavLM-Large [16], which are effective for phoneme recognition due to their rich linguistic content. However, features from these higher layers degraded speaker identity and prosody reproduction, both essential for multi-speaker TTS [29]. Based on this observation, we adopt layer 6 of WavLM-Large, following kNN-VC, as it better preserves speaker-specific information.
@@ IV-A1 Speech-unit-tokenizer
The WavLM-Large encoder produces continuous feature vectors for every 20ms of 16kHz of audio. We discretize these features using $k$-means clustering, as described in Section III-A1. For our proposed method, we use 2000 cluster centers. This choice was made after initial experiments with 100 and 500 clusters, where we observed that more fine-grained quantization led to improved intelligibility of the synthesized speech.
@@ IV-A2 Text-unit-tokenizer
We modify the FastSpeech2 architecture [27] to design a text-to-semantic-unit predictor, replacing Mel spectrogram prediction with semantic unit prediction. The ground truth units are derived by applying the SpeechTokenizer on the ground truth audio. The network is optimized using the Adam optimizer with a learning rate of $45e-$ 4 and a batch size corresponding to 10,000 phonemes. Training is performed on an NVIDIA RTX 3090 GPU, where the network converges within approximately 20 minutes, or after around 10,000 steps. Our implementation builds upon the publicly available SpeechLM framework [30].
3) Vocoder: We use a HiFi-GAN V1 architecture from [31]. We start with the pre-trained vocoder model from the kNN-VC implementation [^1] https://github.com/bshall/knn-vc. We fine-tune two vocoders with frames selected by proposed algorithms: one where frame selection is based solely on inverse k-means sampling, and another where frame selection incorporates both inverse k-means sampling and sub-sequence matching.
IV-B Baselines
We use three SOTA multi-speaker TTS baselines to compare the performance of our proposed framework: the official releases of YourTTS [^2] https://github.com/Edresson/YourTTS [7] and XTTS-v2333https://huggingface.co/coqui/XTTS-v2 [3] with default parameters. We use Amphion’s VALL-E  [^4] https://github.com/open-mmlab/Amphion/tree/main/egs/tts/VALLE_V2 [2] implementation.
IV-C Training and test data
We utilize LibriSpeech train-clean-100 dataset [32] for training both the text-to-semantic-unit tokenizer and the HiFi-GAN vocoder. This dataset comprises 100 hours of text-audio pairs, sampled at 16kHz, and features 251 speakers. We use the LibriSpeech dev-clean subset for validation and LibriTTS-R test-clean [33] for testing. For the LibriTTS-R test-clean set, we select speakers who were not seen during training and have at least 20 utterances, each containing between 3 and 30 words. This results in 31 eligible speakers and a total of 620 utterances. For the VCTK [34] subset which is unseen during training, we sampled 60 speakers where each speaker utters 10 randomly selected sentences, resulting in a total of 600 utterances. We use 5 minutes of reference audio for all variants of SelectTTS, as well as for YourTTS and XTTS-v2, across all evaluation settings. For VALL-E, due to computational constraints, the reference speech was limited to 10-15 seconds.
TABLE I: Objective evaluation results (LibriTTS-R test-clean)